OpenStudy (anonymous):

What is the standard form of (1+2sqrt-3)/(1+sqrt-3)?

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }+\frac{ \sqrt{3} }{ 4 }i\]

OpenStudy (anonymous):

im not that confident in that answer but tell me if it was right or not

OpenStudy (anonymous):

How did you get the answer?

OpenStudy (anonymous):

times it with -sqrt 1 on both sides of the equation

OpenStudy (anonymous):

Okay. Thank you.

OpenStudy (anonymous):

welcome